OpenStudy (anonymous):

can someone explain me how to factor this with the steps plase: 12x^2-15x+3

OpenStudy (unklerhaukus):

all terms have a factor of three, factor that out \[12x^2-15x+3\]\[=3(4x^2-5x+1)\]

OpenStudy (anonymous):

yep... then factor tat trinomial...

OpenStudy (anonymous):

**that

OpenStudy (unklerhaukus):

\[=3(4x^2−5x+1)\]\[=3(4x-1)(x-1)\]

OpenStudy (anonymous):

thank you both :)